# Persistent and anti-Persistent Motion in Bounded and Unbounded Space: Resolution of the First-Passage Problem

### v1

@@ Appendix B The Structure Function
When $f=1$, in the periodically bounded domain, one can show that the walker will be at $n_{0}$ with certainty every $t=\mathfrak{c}N$ steps ($\mathfrak{c}\in\mathbb{N}\cup\{0\}$), that is the system is a Markov chain with periodicity $N$. To do so, let us consider the one-dimensional case (the analysis is readily extended to arbitrary dimensions), where $\bm{\lambda}\left(\frac{2\pi\kappa}{N}\right)=\left(\begin{smallmatrix}e^{\frac{2\pi i\kappa}{N}}&0\\ 0&e^{\frac{-2\pi i\kappa}{N}}\end{smallmatrix}\right)$. Consequently, $\bm{P}_{n_{0}}^{(p)}(n,t)=N^{-1}\sum_{\kappa=0}^{N-1}\exp\left(-\frac{2\pi i\kappa(n-n_{0})}{N}\right)\left(\begin{smallmatrix}e^{\frac{2\pi i\kappa t}{N}}&0\\ 0&e^{\frac{-2\pi i\kappa t}{N}}\end{smallmatrix}\right)\cdot\left[\begin{matrix}\alpha_{1}\\ \alpha_{2}\end{matrix}\right]$. Upon inspection, one clearly sees that when $t=\mathfrak{c}N$ and $n=n_{0}$, the propagator reduces to $\bm{P}_{n_{0}}^{(p)}(n,t)=\left[\begin{matrix}\alpha_{1}&\alpha_{2}\end{matrix}\right]^{\intercal}$, such that $P^{(p)}_{n_{0}}(n_{0},\mathfrak{c}N)=1$.
Appendix C Defect Technique in Correlated Random Walks
We consider the (unordered) set $\mathdutchcal{S}=\left\{(\bm{s}_{i},m_{\bm{s}_{i}}),...,(\bm{s}_{S},m_{\bm{s}_{S}})\right\}$ of $S$ localised targets. To begin, we write, in arbitrary dimensions and with arbitrary numbers of internal degrees of freedom, the Master equation for the state level dynamics
$P(\bm{n},m,t+1)=\sum_{\bm{n}^{\prime}}\sum_{m^{\prime}=1}^{M}$ $\bigg{[}A(\bm{n},m,\bm{n}^{\prime},m^{\prime})P(\bm{n}^{\prime},m^{\prime},t)+\sum_{i=1}^{S}\delta_{\bm{n},\bm{s}_{i}}\delta_{m,m_{\bm{s}_{i}}}(1-\rho_{\bm{s}_{i},m_{\bm{s}_{i}}})A(\bm{s}_{i},m_{\bm{s}_{i}},\bm{n}^{\prime},m^{\prime})P(\bm{n}^{\prime},m^{\prime},t)\bigg{]},$ (30)
where $A(\bm{n},m,\bm{n^{\prime}},m^{\prime})$ is the transition probability from state $\bm{n}^{\prime},m^{\prime}$ to state $\bm{n},m$, and $\rho_{\bm{b}_{i},m_{\bm{b}_{i}}}$, $(0<\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\leq 1)$ governs the probability of getting absorbed at defect $\bm{b}_{i},m_{\bm{b}_{i}}$ where $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}=1$ represents perfect trapping efficiency at that site. To proceed, one first considers the $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\neq 1$ case.
$\widetilde{P}^{(a)}(\bm{n},m,z)$ $=\!\sum_{\bm{n}_{0}^{\prime}}\sum_{m_{0}^{\prime}=1}^{M}\!\bigg{[}\widetilde{P}_{\bm{n}_{0}^{\prime},m_{0}^{\prime}}(\bm{n},m,z)P(\bm{n},m,0)\!+\!\sum_{i=1}^{S}\frac{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}}{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\!-\!1}\widetilde{P}_{\bm{s}_{i},m_{\bm{s}_{i}}}(\bm{n},m,\!z)\!\left[\widetilde{P}^{(a)}(\bm{s},m_{\bm{s}_{i}},\!z)\!-\!P^{(a)}(\bm{s},m_{\bm{s}_{i}},0)\right]\!\bigg{]}\!.$ (32)
where we have used $\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)$ to denote any valid (see the main text) defect free propagator.
$\widetilde{P}^{(a)}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)=\alpha_{m_{0}}\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)+\sum_{i=1}^{S}\frac{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}}{\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}-1}\alpha_{m_{\bm{s}_{i}}}\widetilde{P}_{\bm{s}_{i},m_{\bm{s}_{i}}}(\bm{n},m,z)\widetilde{P}^{(a)}_{\bm{n}_{0},m_{0}}(\bm{s},m_{\bm{s}_{i}},z).$ (33)
and following the standard defect technique procedure (see the Supplementary Material [41] for further discussion) i.e., solving via Cramer’s rule [63] and taking $\rho_{\bm{s}_{i},m_{\bm{s}_{i}}}\to 1$ for all $(\bm{s}_{i},m_{\bm{s}_{i}})$, we find
$\widetilde{P}^{(a)}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)=\alpha_{m_{0}}\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)-\sum_{i=1}^{S}\alpha_{m_{\bm{s}_{i}}}\widetilde{P}_{\bm{s},m_{\bm{s}_{i}}}(\bm{n},m,z)\frac{\det[\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)]}{\det[\mathbb{H}(z)]},$ (34)
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(j)}(\bm{n}_{0},m_{0},z)$ is the same but with the $j^{\text{th}}$ column replaced with $\alpha_{m_{0}}\left[\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{s}_{S},m_{\bm{s}_{S}},z)\right]^{\intercal}$, i.e, they are all known propagators via $\alpha_{m_{0}}\widetilde{P}_{\bm{n}_{0},m_{0}}(\bm{n},m,z)=\alpha_{m_{0}}\bm{e}_{m}^{\intercal}\cdot\bm{\widetilde{P}}_{\bm{n}_{0}}(\bm{n},z)\cdot\bm{e}_{m_{0}}$.
Since $\widetilde{P}^{(a)}_{\bm{n}_{0}}(\bm{n},z)=\sum_{m=1}^{M}\sum_{j=1}^{M}\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}^{(a)}(\bm{n},m,z)$, and $\widetilde{P}^{(\gamma)}_{\bm{n}_{0}}(\bm{n},z)=\sum_{m=1}^{M}\sum_{j=1}^{M}\alpha_{m_{0_{j}}}\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}^{(\gamma)}(\bm{n},m,z)$, where there is no $\alpha_{m_{0}}$ multiplier on the absorbing case as it is already implemented via the initial condition in Eq. (33), one may perform the double summation over both sides of Eq. (34) to arrive at Eq. (21) of the main text.
@@ Appendix C Defect Technique in Correlated Random Walks
$f=\alpha-\frac{\Delta t}{2T},\quad b=\frac{\Delta t}{2T}\quad\text{and}\quad c=1-\alpha$ (S2)
$0\leq\alpha\leq 1$ where $T$ has the dimensions of time and $\lim_{\Delta x,\Delta t\to 0}\frac{\Delta x}{\Delta t}=v$. After some algebra, dividing through by $\Delta t$ and letting $\Delta t$ and $\Delta x$ tend to zero, we find
@@
where the structure of the matrix $\bm{\lambda}(\xi_{1},...,\xi_{d})$ is discussed in Appendix B of the main text. As noted in ref. [42], to obtain the state level probability one needs only to integrate over the specific element in the matrix in Eq. (S11), e.g., $\widetilde{Q}_{\bm{n}_{0},m_{0}}(\bm{n}_{1},...,\bm{n}_{d},m,z)=(2\pi)^{-d}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\bm{e}_{m}^{\intercal}\cdot\widetilde{\bm{Q}}_{\bm{n}_{0}}(\bm{n}_{1},...,\bm{n}_{d},z)\cdot\bm{e}_{m_{0}}\text{d}^{d}\xi$.
@@
$\begin{bmatrix}\widetilde{Q}_{n_{0},1}(n,1,z)\\ \widetilde{Q}_{n_{0},1}(n,2,z)\end{bmatrix}=\frac{\alpha_{1}}{(1-2cz)\sqrt{\left(1+\delta z^{2}\right)^{2}-4\varepsilon^{2}z^{2}}}\begin{bmatrix}\frac{1-cz}{r_{+}(z)^{|n-n_{0}|}}-\frac{az}{r_{+}(z)^{|n-n_{0}+1|}}\\ \frac{bz}{r_{+}(z)^{|n-n_{0}+1|}}\end{bmatrix}$ (S15)
@@
$\begin{bmatrix}\widetilde{Q}_{n_{0},2}(n,1,z)\\ \widetilde{Q}_{n_{0},2}(n,2,z)\end{bmatrix}=\frac{\alpha_{2}}{(1-2cz)\sqrt{\left(1+\delta z^{2}\right)^{2}-4\varepsilon^{2}z^{2}}}\begin{bmatrix}\frac{bz}{r_{+}(z)^{|n-n_{0}-1|}}\\ \frac{1-cz}{r_{+}(z)^{|n-n_{0}|}}-\frac{az}{r_{+}(z)^{|n-n_{0}-1|}},\end{bmatrix}.$ (S16)
@@
$\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}\mu_{d}\delta(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu})=\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}e^{-i\bm{\kappa}^{\intercal}\cdot\bm{N}\bm{\xi}}$ (S20)
where $\mu_{d}=\frac{2\pi}{N_{d}}$, one finds
$\widetilde{\bm{P}}^{(p)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)=\frac{1}{N^{d}}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}e^{-i(\bm{n}^{\intercal}-\bm{n_{0}}^{\intercal})\cdot\bm{\xi}}\delta(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu})\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},...,\xi_{d})\right]^{-1}\cdot\bm{U_{m_{0}}}\text{d}^{d}\xi.$ (S21)
Since the region of integration is over the Brillouin zone of the lattice sites, the singularities of $\delta(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu})$ lie in the integral region if and only if $\bm{\kappa}$ is a site in the lattice (for a general discussion of solving integrals of this type see [42]). Solving the integral in Eq. (S21) using this property leads straightforwardly to
@@
which after taking the inverse $z$-transform via the identity $\sum_{t=0}^{\infty}z^{t}\bm{\lambda}(\bm{\xi})^{t}=\left[\mathbb{I}-z\bm{\lambda}(\bm{\xi})\right]$, one finds Eq. (14) of the main text.

### v2

@@ Appendix B The Structure Function
When $f=1$, in the periodically bounded domain, one can show that the walker will be at $n_{0}$ with certainty every $t=\mathfrak{c}N$ steps ($\mathfrak{c}\in\mathbb{N}\cup\{0\}$), that is the system is a Markov chain with periodicity $N$. To do so, let us consider the one-dimensional case (the analysis is readily extended to arbitrary dimensions), where $\bm{\lambda}\left(\frac{2\pi\kappa}{N}\right)=\left(\begin{smallmatrix}e^{\frac{2\pi i\kappa}{N}}&0\\ 0&e^{\frac{-2\pi i\kappa}{N}}\end{smallmatrix}\right)$. Consequently, $\bm{P}_{n_{0}}^{(p)}(n,t)=N^{-1}\sum_{\kappa=0}^{N-1}\exp\left(-\frac{2\pi i\kappa(n-n_{0})}{N}\right)\left(\begin{smallmatrix}e^{\frac{2\pi i\kappa t}{N}}&0\\ 0&e^{\frac{-2\pi i\kappa t}{N}}\end{smallmatrix}\right)\cdot\left[\begin{matrix}\alpha_{1}\\ \alpha_{2}\end{matrix}\right]$. Upon inspection, one clearly sees that when $t=\mathfrak{c}N$ and $n=n_{0}$, the propagator reduces to $\bm{P}_{n_{0}}^{(p)}(n,t)=\left[\begin{matrix}\alpha_{1}\\ \alpha_{2}\end{matrix}\right]$, such that $P^{(p)}_{n_{0}}(n_{0},\mathfrak{c}N)=1$.
Appendix C The Hexagonal Propagator
With $Z=6$, the governing equations for correlated motion on the hexagonal lattice comprisies of six coupled equations, which we write explicitly in the Supplementary Material (ref. [45]). With the initial condition $\bm{U}_{m_{0}}=\delta_{n_{1},n_{0_{1}}}\delta_{n_{2},n_{0_{2}}}[\alpha_{1},...,\alpha_{6}]^{\intercal}$, and following the same procedure as for the hypercubic lattice (Sec. Derivation of the Unbounded Propagator of ref. [45]), it is possible to write the lattice Green’s function for the hexagonal lattice as
$\widetilde{\bm{Q}}^{(\mathcal{H})}_{n_{0}}(\bm{n},z)=\frac{1}{4\pi^{2}}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}e^{-i\bm{\xi}(\bm{n}-\bm{n}_{0})}\bm{\lambda}^{(\mathcal{H})}(\xi_{1},\xi_{2})\cdot\bm{U_{m_{0}}}\text{d}\xi_{1}\text{d}\xi_{2},$ (37)
where
$\bm{\lambda}^{(\mathcal{H})}(\xi_{1},\xi_{2})=$ (38)
[…]
with the elements representing the different movement options are labeled, in Eq. (38), via the arrows. For clarity we also show pictorially some movement options in Fig. 12.
Figure 12: (Colour Online). A schematic representing two of the six permissible movement steps in the correlated hexagonal random walk where the solid arrows represent the movement probabilities if the walker had entered the site via the dotted arrow. The movements shown on the left side case are shown in the $3^{\text{rd}}$ column of Eq. (38) of the main text and the movements shows in the right panel are encoded in the $2^{\text{nd}}$ column of Eq. (38) of the main text.
$\widetilde{\bm{P}}^{(\mathcal{H})}_{\bm{n}_{0}}(n_{1},n_{2},z)=\sum_{\kappa_{1}=-\infty}^{\infty}\sum_{\kappa_{2}=-\infty}^{\infty}\widetilde{\bm{Q}}_{\bm{n}_{0}}^{(\mathcal{H})}(n_{1}+\hat{n}_{1},n_{2}+\hat{n}_{2},z),$ (39)
where
$\begin{bmatrix}\widehat{n}_{1}\\ \widehat{n}_{2}\end{bmatrix}=\begin{bmatrix}-R\kappa_{1}+(2R+1)\kappa_{2}\\ -(R+1)\kappa_{1}-R\kappa_{2}\end{bmatrix},$ (40)
and since, like the hypercubic case, the periodic boundary conditions for the correlated random walk are identical to their diffusive counterpart, Eq.  (37) can be solved explicitly following the same procedure as the solution to the LRW with periodic boundary conditions derived in [59]. This leads to Eq. (20) of the main text.
As mentioned in Sec. IV.2 of the main text, the non-orthogonality of the coordinates leads to a ‘shift’ in the position of the walker when traversing the periodic boundary. For a deeper discussion we again refer the interested reader to [59], and provide here a simple explanation of the dynamics seen in Fig. 4 of the main text by describing the path taken by the probability pulse that is near North-East of the domain in the ballistic case depicted in Fig. 4(a) (other pulses may then be understood similarly due to the rotational symmetry of the system).
As mentioned in Sec. IV.2 of the main text, the shift in the boundary condition means that the walker will hit every site in the finite domain with certainty even if $f=1$, something that is not the case in the $d>1$ hypercubic lattice. We show this formally here by considering the $z$-transfrom of Eq. (20) may be written in the form [46]
$\widetilde{\bm{P}}^{(\mathcal{H})}_{\bm{n}_{0}}(\bm{n},z)=\frac{1}{\Omega}\left[\mathbb{I}-z\bm{\lambda}^{(\mathcal{H})}(0,0)\right]^{-1}+\widetilde{\bm{P}}^{(\kappa\neq 0)}_{\bm{n}_{0}}(\bm{n},z),$ (41)
where $\widetilde{P}^{(\kappa\neq 0)}_{\bm{n}_{0}}(\bm{n},z)$ is the generating function of the double summation in Eq. (20). Since $\bm{\lambda}^{(\mathcal{H})}(0,0)=\mathbb{I}$ for $(\kappa_{1},\kappa_{2})=(0,0)$ only, it is clear that at $z=1$, each element $\widetilde{\bm{P}}^{(\kappa\neq 0)}_{\bm{n}_{0}}(\bm{n},z)$ is holomorphic. Therefore, $\widetilde{\bm{P}}^{(\mathcal{H})}_{\bm{n}_{0}}(\bm{n},z)\to\infty$ as $z\to 1$ showing it is certain that all sites will be hit eventually. This is true for all movement parameters except when $b=1$ or $\ell_{i}=1$.
Appendix D Alternative Derivation of the Single Target First-Passage Probability
We generalise here the approach taken in ref. [34]. To do so, we allow for an arbitrary initial weighting distribution and make no assumptions about the symmetry of the system meaning it is valid for any boundary condition.
For simplicity in the algebra we assume without loss of generality that the system is one-dimensional with two internal states $m=\{1,2\}$. The equations relating the state-level occupation probability and state-level first passage are written as:
$\alpha_{m_{0_{1}}}P_{n_{0},1}(s,1,t)$ $=\alpha_{m_{0_{1}}}\delta_{m_{0},1}\delta_{t,0}\delta_{n,n_{0}}$ (42)
$\quad\quad\quad+\sum_{t^{\prime}=0}^{t}F_{n_{0},1}(s,1,t^{\prime})P_{s,1}(s,1,t-t^{\prime})+F_{n_{0},1}(s,2,t^{\prime})P_{s,2}(s,1,t-t^{\prime}),$
$\alpha_{m_{0_{2}}}P_{n_{0},2}(s,1,t)$ $=\alpha_{m_{0_{2}}}\delta_{m_{0},1}\delta_{t,0}\delta_{n,n_{0}}$
$\quad\quad\quad+\sum_{t^{\prime}=0}^{t}F_{n_{0},2}(s,1,t^{\prime})P_{s,1}(s,1,t-t^{\prime})+F_{n_{0},2}(s,2,t^{\prime})P_{s,2}(s,1,t-t^{\prime}),$
$\alpha_{m_{0_{1}}}P_{n_{0},1}(s,2,t)$ $=\alpha_{m_{0_{1}}}\delta_{m_{0},2}\delta_{t,0}\delta_{n,n_{0}}$
$\quad\quad\quad+\sum_{t^{\prime}=0}^{t}F_{n_{0},1}(s,1,t^{\prime})P_{s,1}(s,2,t-t^{\prime})+F_{n_{0},1}(s,2,t^{\prime})P_{s,2}(s,2,t-t^{\prime}),$
$\alpha_{m_{0_{2}}}P_{n_{0},2}(s,2,t)$ $=\alpha_{m_{0_{2}}}\delta_{m_{0},2}\delta_{t,0}\delta_{n,n_{0}}$
$\quad\quad\quad+\sum_{t^{\prime}=0}^{t}F_{n_{0},2}(s,1,t^{\prime})P_{s,1}(s,2,t-t^{\prime})+F_{n_{0},2}(s,2,t^{\prime})P_{s,2}(s,2,t-t^{\prime}).$
Equation (42) reduces to Eq. (2) of ref. [34] with the choice of $\alpha_{m_{0_{1}}}=1$, $\alpha_{m_{0_{2}}}=0$, which implies $F_{n_{0},2}(s,1,t)=F_{n_{0},2}(s,2,t)=0$ as no trajectories start in state 2.
Proceeding in the $z$-domain, and assuming $s\neq n_{0}$, we find
$\alpha_{m_{0_{1}}}\widetilde{P}_{n_{0},1}(s,1,z)$ $=\widetilde{F}_{n_{0},1}(s,1,z)\widetilde{P}_{s,1}(s,1,z)+\widetilde{F}_{n_{0},1}(s,2,z)\widetilde{P}_{s,2}(s,1,z),$ (43)
which one may formulate as the matrix equation
$\!\begin{bmatrix}\widetilde{F}_{n_{0},1}(s,1,z)\\ \widetilde{F}_{n_{0},1}(s,2,z)\\ \widetilde{F}_{n_{0},2}(s,1,z)\\ \widetilde{F}_{n_{0},2}(s,2,z)\end{bmatrix}\!=\!\begin{bmatrix}\widetilde{P}_{s,1}(s,1,z)&\widetilde{P}_{s,2}(s,1,z)&0&0\\ 0&0&\widetilde{P}_{s,1}(s,1,z)&\widetilde{P}_{s,2}(s,1,z)\\ \widetilde{P}_{s,1}(s,2,z)&\widetilde{P}_{s,2}(s,2,z)&0&0\\ 0&0&\widetilde{P}_{s,1}(s,2,z)&\widetilde{P}_{s,2}(s,2,z)\end{bmatrix}^{-1}\!\cdot\!\begin{bmatrix}\alpha_{m_{0_{1}}}\widetilde{P}_{n_{0},1}(s,1,z)\\ \alpha_{m_{0_{2}}}\widetilde{P}_{n_{0},2}(s,1,z)\\ \alpha_{m_{0_{1}}}\widetilde{P}_{n_{0},1}(s,2,z)\\ \alpha_{m_{0_{2}}}\widetilde{P}_{n_{0},2}(s,2,z)\end{bmatrix}\!.$ (44)
Upon evaluating Eq. (44), one finds Eq. (30) of the main text by performing $\sum_{m_{0}=1}^{2}\sum_{m=1}^{2}\widetilde{F}_{n_{0},m_{0}}(s,m,z)$, namely performing a summation over the column vector of state level first-passages.
The procedure detailed here generalises trivially to a target of arbitrary number of states (i.e., it is valid in any dimension), however the matrix that requires inversion may become large. In such cases, the approach via defective sites presented in the main text is more suitable.
Pertinent to the directional first-passage as well as the first-passage in the unbounded one-dimensional case (as the directional first-passage equals the first-passage in this case) is the situation of one target state at some $\bm{s},m_{\bm{s}}$. We focus on the arbitrarily dimensional directional first-passage probability. In such a situation, the renewal equations simplify to
$\alpha_{m_{0_{1}}}P_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},t)$ $=\alpha_{m_{0_{1}}}\delta_{m_{0},1}\delta_{t,0}\delta_{\bm{s},\bm{n}_{0}}+\sum_{t^{\prime}=0}^{t}F_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},t^{\prime})P_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},t-t^{\prime}),$ (45)
[…]
We again assume $\bm{s}\neq\bm{n}_{0}$ and take the generating function to find
$\alpha_{m_{0_{1}}}\widetilde{P}_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},z)$ $=\widetilde{F}_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},z)\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z),$ (46)
[…]
For each equation in Eq. (46) one may solve for
$\widetilde{F}_{\bm{n}_{0},i}(\bm{s},m_{\bm{s}},z)=\frac{\alpha_{m_{0_{i}}}\widetilde{P}_{\bm{n}_{0},i}(\bm{s},m_{\bm{s}},z)}{\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z)},$ (47)
independently of the other equations and upon summation $\widetilde{F}_{\bm{n}_{0}}(\bm{s},m_{\bm{s}},z)=\sum_{i=1}^{M}\widetilde{F}_{\bm{n}_{0},i}(\bm{s},m_{\bm{s}},z)$, we find Eq. (29) of the main text. By again setting $\alpha_{m_{0_{i}}}=1$ and all other $\alpha_{m_{0_{j}}}=0$ ($j\neq i$) we recover the known cases of refs. [34] and [35].
For specific cases the first-passage expression, Eq. (27), may actually be simplified further. This occurs when $\widetilde{P}_{n,i}(n,i,z)=\widetilde{P}_{n,j}(n,j,z)$ and $\widetilde{P}_{n,j}(n,i,z)=\widetilde{P}_{n,i}(n,j,z)$, which is true only when the walk is translationally invariant and we demonstrate this explicitly here in the one-dimensional case.
Inserting the above identities into Eq. (30) we find
$\widetilde{F}_{n_{0}}(s,z)=$
$\frac{\left\{\!\alpha_{1}\!\left[\!\widetilde{P}_{n_{0},1}(s,1,z)\!+\!\widetilde{P}_{n_{0},1}(s,2,z)\right]\!+\!\alpha_{2}\left[\widetilde{P}_{n_{0},2}(s,1,z)\!+\!\widetilde{P}_{n_{0},2}(s,2,z)\!\right]\right\}\!\left[\!\widetilde{P}_{s,2}(s,2,z)\!-\!\widetilde{P}_{s,1}(s,2,z)\!\right]}{\widetilde{P}_{s,2}(s,2,z)^{2}-\widetilde{P}_{s,1}(s,2,z)^{2}}\!.$ (48)
Since, $\alpha_{1}\left[\widetilde{P}_{n_{0},1}(s,1,z)+\widetilde{P}_{n_{0},1}(s,2,z)\right]+\alpha_{2}\left[\widetilde{P}_{n_{0},2}(s,1,z)+\widetilde{P}_{n_{0},2}(s,2,z)\right]=\widetilde{P}_{n_{0}}(s,z)$, Eq. (48) may be rewritten as
$\widetilde{F}_{n_{0}}(s,z)=\frac{\widetilde{P}_{n_{0}}(s,z)\left[\widetilde{P}_{s,2}(s,2,z)-\widetilde{P}_{s,1}(s,2,z)\right]}{\widetilde{P}_{s,2}(s,2,z)^{2}-\widetilde{P}_{s,1}(s,2,z)^{2}},$ (49)
and then upon using the differences of two squares property on the denominator we find
$\widetilde{F}_{n_{0}}(s,z)=\frac{\widetilde{P}_{n_{0}}(s,z)}{\widetilde{P}_{s,2}(s,2,z)+\widetilde{P}_{s,1}(s,2,z)}.$ (50)
To perform the final simplification we note that with $\alpha_{1}+\alpha_{2}=1$, the denominator may be rewritten as $(\alpha_{1}+\alpha_{2})\left[\widetilde{P}_{s,2}(s,2,z)+\widetilde{P}_{s,1}(s,2,z)\right]$. By multiplying this expression out and once again using the properties of the translationally invariant propagators given above, we see that
$(\alpha_{1}+\alpha_{2})$ $\left[\widetilde{P}_{s,2}(s,2,z)+\widetilde{P}_{s,1}(s,2,z)\right]=$
$\alpha_{1}\left[\widetilde{P}_{s,1}(s,1,z)+\widetilde{P}_{s,1}(s,2,z)\right]+\alpha_{2}\left[\widetilde{P}_{s,2}(s,1,z)+\widetilde{P}_{s,2}(s,2,z)\right],$ (51)
where it is clear from the LHS of Eq. (51) that we may re-write the denominator of Eq. (50) as $\widetilde{P}_{s}(s,z)$ leading to
$\widetilde{F}_{n_{0}}(s,z)=\frac{\widetilde{P}_{n_{0}}(s,z)}{\widetilde{P}_{s}(s,z)}.$ (52)
@@ Appendix D Alternative Derivation of the Single Target First-Passage Probability
$f=\alpha-\frac{\Delta t}{2T},\quad b=\frac{\Delta t}{2T}\quad\text{and}\quad c=1-\alpha,$ (S2)
where $0\leq\alpha\leq 1$, $T$ has the dimensions of time and $\lim_{\Delta x,\Delta t\to 0}\frac{\Delta x}{\Delta t}=v$. After some algebra, dividing through by $\Delta t$ and letting $\Delta t$ and $\Delta x$ tend to zero, we find
@@
where the structure of the matrix $\bm{\lambda}(\xi_{1},...,\xi_{d})$ is discussed in B of the main text. As noted in ref. [46], to obtain the state level probability one needs only to integrate over the specific element in the matrix in Eq. (S28), e.g., $\widetilde{Q}_{\bm{n}_{0},m_{0}}(\bm{n}_{1},...,\bm{n}_{d},m,z)=(2\pi)^{-d}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\bm{e}_{m}^{\intercal}\cdot\widetilde{\bm{Q}}_{\bm{n}_{0}}(\bm{n}_{1},...,\bm{n}_{d},z)\cdot\bm{e}_{m_{0}}\text{d}^{d}\xi$.
@@
$\begin{bmatrix}\widetilde{Q}_{n_{0},1}(n,1,z)\\ \widetilde{Q}_{n_{0},1}(n,2,z)\end{bmatrix}=\frac{\alpha_{1}}{(1-2cz)\sqrt{\left(1+\delta z^{2}\right)^{2}-4\varepsilon^{2}z^{2}}}\begin{bmatrix}\frac{1-cz}{r_{+}(z)^{|n-n_{0}|}}-\frac{fz}{r_{+}(z)^{|n-n_{0}+1|}}\\ \frac{bz}{r_{+}(z)^{|n-n_{0}+1|}}\end{bmatrix}$ (S15)
@@
$\begin{bmatrix}\widetilde{Q}_{n_{0},2}(n,1,z)\\ \widetilde{Q}_{n_{0},2}(n,2,z)\end{bmatrix}=\frac{\alpha_{2}}{(1-2cz)\sqrt{\left(1+\delta z^{2}\right)^{2}-4\varepsilon^{2}z^{2}}}\begin{bmatrix}\frac{bz}{r_{+}(z)^{|n-n_{0}-1|}}\\ \frac{1-cz}{r_{+}(z)^{|n-n_{0}|}}-\frac{fz}{r_{+}(z)^{|n-n_{0}-1|}},\end{bmatrix}.$ (S16)
@@
$\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}\mu^{d}\delta(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu})=\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}e^{-i\bm{\kappa}^{\intercal}\cdot\bm{N}\bm{\xi}}$ (S20)
where we define $\mu^{d}=\prod_{j=1}^{d}2\pi/N_{j}$ and $\bm{\mu}=\left[2\pi/N_{1},...,2\pi/N_{d}\right]$, one finds
$\widetilde{\bm{P}}^{(p)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)\!=\!\frac{1}{N^{d}}\int_{-\varepsilon_{1}}^{2\pi-\varepsilon_{1}}...\int_{-\varepsilon_{d}}^{2\pi-\varepsilon_{d}}\!\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}\!e^{-i(\bm{n}^{\intercal}-\bm{n_{0}}^{\intercal})\cdot\bm{\xi}}\delta(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu})\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},...,\xi_{d})\right]^{-1}\cdot\bm{U_{m_{0}}}\text{d}^{d}\xi,$ (S21)
where $N^{d}=\prod_{j=1}^{d}N_{j}$ and the periodicity of the function has allowed us to replace the region of integration from $V=[-\pi,\pi]^{d}$ to $[-\varepsilon_{d},2\pi-\varepsilon_{d}]$ along each dimension, where $0<\varepsilon_{d}<2\pi/N_{d}$. The singularities of $\delta(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu})$ lie in the integral region if and only if $\bm{\kappa}$ is a site in the lattice (for a general discussion of solving integrals of this type see [46]). We may use this property to perform the intergrals in Eq. (S21) and obtain
@@
which after taking the inverse $z$-transform via the identity $\sum_{t=0}^{\infty}z^{t}\bm{\lambda}(\bm{\xi})^{t}=\left[\mathbb{I}-z\bm{\lambda}(\bm{\xi})\right]^{-1}$, one finds Eq. (14) of the main text.